# Unpacking Personal(?!) Health Informatics: An Investigation of Awareness, Understanding, & Leveraged Utility in India

### v1

@@ 5. Discussion
Our findings shed light on how urban Indians adopt, use, and perceive health wearables for PHI in India, emphasizing the cultural, societal, and technological factors influencing user behavior. While existing literature has focused on specialized applications of wearables, such as fall detection or acute event alerts (Delahoz and Labrador, 2014; Gjoreski et al., 2016; Bat-Erdene and Saver, 2021; Amato et al., 2022; Ghosh and Das, 2024), we uncover a broader gap: how these same technologies are—or could be—woven into routine health management practices. By situating our findings within both ecological systems theory (Bronfenbrenner, 1974) and Wozniak et al.’s concept of Health Information Ecologies (Wozniak-O’Connor et al., 2024), we illustrate how India’s collectivist familial networks, cultural expectations, and evolving technology landscapes jointly shape PHI adoption. Below, we discuss three key themes—navigating a fragmented ecosystem, wearables as care devices, and culturally sensitive engagement —and offer design and policy recommendations to support a more culturally attuned PHI future.
5.1. Navigating a Fragmented Digital Health Ecosystem: Technology & Policy
India’s digital health landscape is marked by fragmentation and the absence of centralized, reliable information—conditions that contrast sharply with the integrated systems typical in Western contexts. This fragmentation hampers PHI usability, trust, and effectiveness.
Participants reported difficulty navigating multiple disjointed platforms, often defaulting to traditional health practices. Choosing a tracker was perceived as a significant investment, complicated by non-compatible ecosystems—a clear barrier (section 4.2). While PHI systems evolve to support personal health management, the IT infrastructure backing them remains outdated and fragmented (Jones et al., 2012). Proprietary ecosystems hinder interoperability, complicating and increasing the cost of decision-making. For example, a Samsung smartwatch user with an iPhone or non-Samsung Android device cannot access all features or rely on data accuracy, illustrating how such silos generate user confusion and disengagement. This calls for design strategies that reduce the cognitive load of choice and simplify use. As seen in our prototype evaluation, users need flexible interaction modes and wearable form factors tailored to their comfort levels (section 4.4.2). While some of these challenges can be overcome through software/interaction design, the rest require policy-level changes that enforce manufacturers to comply with an open data framework that ensures accessibility of health data regardless of the medium or proprietary ecosystem requirements of individual brands/companies.
Furthermore, the fragmented digital ecosystem and the accompanying mistrust raise ethical and privacy concerns. This aspect of PHI must be balanced with the need for integrated, actionable health information. While countries worldwide have policies regarding collecting and using health information (for example, HIPAA, PHIPA, GDPR, and APPs), India still uses the IT Act of 2000 even for medical and health information. The IT Act 2000 provides basic data protection by requiring data owners to be informed about its use, but it is not tailored to healthcare. The National Digital Health Blueprint (NDHB) recommends ISO 22600:2014 Health Informatics, but implementing these standards may be challenging for many healthcare providers in India, and ensuring compliance at the provider level could also be challenging for the government. Although similar healthcare-specific legislation was proposed in 2017 and 2022, it has not been implemented. While the Government of India has now published the “Digital Personal Data Protection Act, 2023” (DPDP 2023) covering all personal data (including health data), it remains unenforced and is not robust enough to meet the requirements posed by Indian users’ socially reflective shared caring models.
We propose a unified, open framework that aggregates PHI, doctor visits, lab reports, medications, and more into a single interface. This interface should support all smart wearables—focused on Android, which dominates the Indian market (94%)—and employs accessible language to reduce learning burdens and foster proactive health behaviors.
5.2. Wearables as Care Devices: Collective Care & Engagement Strategies
Our participants overwhelmingly described health wearables not merely as personal gadgets but as instruments of family‐centered care (findings 4.1.1). Unlike the individualist framing common in Western studies (Chmielewski and Meyer, 2024), Indian users envision sharing health metrics/analysis with parents, spouses, or children, turning data streams into collaborative decision‐making tools. We observed that users readily assume caregiver roles—monitoring a parent’s daily steps or a child’s sleep patterns —and express enthusiasm for co‐designed interfaces that let family dyads explore and interpret data together (Lieber et al., 2020; Ugargol and Bailey, 2018; Gustafsson et al., 2022; Bhat et al., 2023).
To drive proactive engagement, gamification of PHI systems must capture activities familiar to Indian lifestyles—such as traditional labor and household tasks—rather than only Western‐style workouts to motivate them better (findings 4.1.2 & 4.3). At the same time, designs should include contextual alerts (for example, flagging sustained deviations in resting heart rate) that prompt timely medical check‐ins, moving users from reactive responses toward ongoing health management. Prior studies on intergenerational sharing of health data (Binda et al., 2017; Cagiltay and Mutlu, 2024; Lee et al., 2024a), women’s data ecosystems (Grimme et al., 2024; Karusala et al., 2023a), and children-specific care for chronic ailments (Cha et al., 2024), reinforce that broadening the scope of shareable metrics can foster more comprehensive, preventive care—particularly where caregiving is a shared, familial endeavor.
Privacy concerns were minimal among our participants (finding 4.1.2), indicating openness to data exchange within trusted circles. This suggests potential for paired or multi‐user wearable configurations, provided that the system offers clear, in-app explanations of how data are collected, processed, and used. Policymakers can foster privacy-conscious behaviors by enforcing clear regulations around data protection, specifically in PHI wearables. As people adopt a more proactive approach, wearables’ role in caregiving could be significantly enhanced by providing continuous insights that allow early intervention before critical incidents. The ability to share comprehensive health data with family members could also strengthen the perceived value of these systems, encouraging consistent use and deeper integration into daily routines. Designers and developers must therefore prioritize features that integrate into users’ lives, facilitating comprehensive health monitoring and sharing. Such usability and transparency, coupled with policy measures mandating robust PHI protections, will be essential to transforming wearables from novelty gadgets into indispensable tools for early intervention and collaborative health management. Our observations showcase that the ”Personal” in PHI involves more than the individual in the Indian context. These family-centered insights directly inform our unified PHI platform vision in the next section.
5.3. Culturally Sensitive Engagement: Norms, Literacy, Misinformation, & Trust
In India, paper-and-analog practices remain deeply trusted, but users embrace digital PHI when it feels like an extension, not a replacement. Participants, for example, described “smart” digital pressure cuffs that look and behave like traditional sphygmomanometers yet auto-sync readings to a unified PHI dashboard (findings 4.4.2). We frame this as a modular hybrid PHI paradigm, where users toggle between digital and analog interfaces based on context and comfort. Such an approach acknowledges India’s healthcare realities and offers a pragmatic, modular pathway for digital transformation—one that integrates seamlessly with generic platforms.
In a society where talking openly about intimate or sexual health is still taboo, cultural norms around sharing vary by metric: intimate health data (e.g., sexual health) was shared only with spouses, whereas vitals like step counts traveled freely among parents, children, or extended family depending on trust (findings 4.1.1). Parents even leaned on devices recommended by their children, illustrating “derived trust.” To embrace these differences, we argue for a culturally informed data-governance model with user-controlled sharing settings and transparent, in-app analytics—so families can collaborate on care without over- or under-sharing. Future policy frameworks should incorporate these principles to foster an ecosystem where digital health innovations are ethically grounded and culturally resonant.
@@ 5.3. Culturally Sensitive Engagement: Norms, Literacy, Misinformation, & Trust
Low health literacy (Bagalkot et al., 2020; Su, 2023; Teixeira Lopes and Ramos, 2020; Egede et al., 2021; Teixeira Lopes and Henriques, 2024; Noblin et al., 2012; Kreps, 2017; Lee et al., 2022; Kemp et al., 2021) and rampant medical misinformation (Pasi, 2024) further compound these challenges and impact adoption. Participants distrusted novel features when metrics seemed “off” or unfamiliar/unevaluated (section 4.2.2), and authoritative Indian health association articles rank poorly in search results from Indian IPs, often supplanted by vague content, whereas Western IP searches surface reliable sources prominently. This disparity highlights the need for authentic, reliable, vetted, user-friendly health information platforms so Indians can access trustworthy guidance without skepticism. And since the pandemic has driven more health-conscious behaviors, implementing the long-overdue 2005 school-health curriculum recommendation (MoHFW, 2005) could foster health-conscious behavior from a young age (Kann et al., 2007; Pulimeno et al., 2020; Auld et al., 2020; Benes and Alperin, 2022; Sulz et al., 2024).
Trust hinges on data accuracy and medical validation. Wearables often function as “black boxes,” limiting lay users’ comprehension (Shin et al., 2019; Zhang and Shahriar, 2020). Sustained adoption, therefore, requires medical validation and clinical integration. Devices should earn formal endorsements, sync with electronic health records, and surface sensor-accuracy indicators (e.g., confidence intervals around SpO2 or heart-rate) to bolster user confidence (Huhn et al., 2022; Volpato et al., 2021; Kang and Exworthy, 2022; Singh et al., 2022a; Samhale, 2022; Zhang et al., 2022; Karnatak et al., 2023; Thapa et al., 2023). Integrating wearables with formal healthcare systems will further bolster their credibility and sustained use.
These collective/shared care insights directly inform our integrated PHI platform that consolidates disparate data, embeds user-controlled sharing and transparent analytics, employs inclusive design methods (e.g., family workshops), and links seamlessly with clinical systems—so that as health literacy grows, users can critically weigh benefits against risks (Mackert et al., 2016; Meyers et al., 2020).
5.4. Future Work
@@
While there are novel recommendations on personal, public, and cultural aspects of PHI adoption, the authors acknowledge several limitations in this study. First, the cross-sectional design captures user behavior and perceptions at a single point in time, limiting the ability to observe how these factors may evolve. Additionally, the focus on urban, middle- and upper-middle-class participants restricts the generalizability of the findings to rural and lower-income populations, where different adoption barriers, such as different lifestyle practices, natural remedies, and many more, may exist. Finally, the reliance on self-reported data also introduces potential bias, as participants may inaccurately report their PHI usage, utility, and adoption.
@@
Building upon our observations and the strategic direction outlined, we propose the design and development of a proof-of-concept application aimed at addressing general health, wellbeing, and medical inquiries. This system will also assist users in interpreting and contextualizing their personal health data aggregated from diverse sources, presenting the information in a format that is both comprehensible and actionable. Leveraging advancements in artificial intelligence, particularly finetuned open-source large language models (LLMs), the application will integrate data from clinical documents (e.g., physician reports, laboratory results, and medication records) alongside information from wearable health devices. This multimodal data synthesis will enable the generation of personalized, holistic insights, thereby enhancing user engagement and health awareness. The interface will employ a conversational, chatbot-style design to facilitate intuitive interaction, mediate family sharing norms, and surface literacy-appropriate explanations. Beyond helping users understand their personal health information (PHI), the system will also provide access to evidence-based health, medical, and wellness content. By sourcing validated information from reputable sources, the application seeks to address the pervasive issue of health misinformation and contribute to the improvement of public health literacy.

### v2

@@ 5. Discussion and Design Implications
Our findings shed light on how urban Indians adopt, use, and perceive PHI for proactive care in India, highlighting the cultural, societal, and technological factors that influence user behavior. While existing literature has focused on specialized applications of wearables, such as fall detection or acute event alerts (Delahoz and Labrador, 2014; Gjoreski et al., 2016; Bat-Erdene and Saver, 2021; Amato et al., 2022; Ghosh and Das, 2024), we identify a broader gap: how these same technologies can be (or could be) integrated into routine health management practices. By situating our findings within both ecological systems theory (Bronfenbrenner, 1974) and Wozniak et al.’s concept of Health Information Ecologies (Wozniak-O’Connor et al., 2024), we illustrate how India’s collectivist familial networks, cultural expectations, and evolving technology landscapes jointly shape the adoption of PHI for proactive care.
Our empirical work answers two linked questions. RQ1 asked how Indians adopt and perceive PHI. We recognize that uptake is shaped by both individual motivations (goal-driven tracking similar to western observations (Chmielewski and Meyer, 2024)) and collective caregiving practices (family-mediated monitoring and context-specific observations reflecting prior work’s findings such as supporting older adults in daily living/patients with chronic conditions/diabetic children with health management and so on (Lieber et al., 2020; Ugargol and Bailey, 2018; Gustafsson et al., 2022; Bhat et al., 2023; Cha et al., 2024)). Cultural norms that privilege embodied cues and clinician confirmation often mean that PHI is engaged with reactively, unless mediated by trusted social or clinical actors. RQ2 inquired about the conditions that promote effective use of PHI. We determine that effective utilization requires (a) agency: granular, transparent control of what is shared and with whom; (b) elicitation: trustworthy signal presentation (confidence badges) and multi-mode prompting strategies; and (c) engagement: literacy-appropriate explanations and family/clinician workflows that translate signals into action. Together, these capabilities form the operational core of PHI-Proact (a conceptual operational map for effective utilization of PHI for proactive care).
5.1. PHI-Proact: Agency, Elicitation, & Engagement
Through a deeper analysis of our findings, we propose PHI-Proact, an operational map that translates ecological PHI insights into prescriptive design levers for proactivity. PHI-Proact foregrounds three interdependent domains:
Agency. Explicit, granular controls for data types, sharing recipients, temporal windows, and role templates (trusted proxies). Agency expands Bronfenbrenner’s (Bronfenbrenner, 1974) actor layer into concrete UI primitives that reflect family and clinician roles in India.
Elicitation. Design patterns for when and how to surface signals: confidence badges, escalation paths (user → family → clinician), scheduled vs context-aware prompts, and clear action pathways. This domain addresses technical trustworthiness and the behavioral mechanics of prompting.
Engagement. Literacy-appropriate explanations, culturally contextual motivators, family dashboards, and clinician-ready summaries that scaffold interpretation and action. This domain operationalizes the “support” layer of ecological models into design patterns.
PHI-Proact therefore extends prior ecological PHI accounts by converting nested socio-technical influences into a concise set of actionable design levers and evaluation metrics suited to non-WEIRD contexts where family dynamics, device heterogeneity, and literacy constraints critically shape PHI adoption.
Our findings confirm established HCI observations (Lazar et al., 2015; Zhang and Shahriar, 2020; Lu et al., 2024) regarding abandonment driven by usability friction and information overload, aligning with prior evidence on the importance of social support and clinician integration for sustained digital health engagement. At the same time, our work diverges from many Western studies that foreground the individual, quantified self-subject. In our Indian sample, family-mediated workflows and system-level interoperability (e.g., EHR compatibility) emerge as primary, context-specific requirements for proactivity. These contrasts suggest that PHI designs should prioritize collective caregiving and policy-level interoperability in non-WEIRD settings.
5.2. Recommendations for Designers and Developers
Based on a critical analysis using our PHI-Proact mapping, we derive ten prescriptive recommendations, each tied to our thematic findings and, where applicable, to prototype validation, and present them below in Table LABEL:Table_5. Each of these recommendations includes one or more of the three aspects of the PHI-Proact mapping. Recommendations R1, R2, R3, and R4 address various aspects of the Agency from the user’s perspective, as well as the stakeholders involved. Recommendations R2, R4, R5, and R8 aim to augment sensing and prompting strategies. Finally, R6, R7, R8, and R9 are focused on revamping engagement and support for all stakeholders involved. Beyond these nine recommendations, utilizing all three aspects of the PHI-Proact mapping, we also provide a final recommendation to ensure that designers and developers employ appropriate evaluation and feedback mechanisms to provide better-aligned support for proactive collective care.
| ††: The table listing recommendations for designers and developers. | |
| R1 | Explicit consent and granular controls: Designs should strike a balance between fine-grained controls and ease of access for the users, allowing users to specify what is shared, who receives it, and for how long, without needing to navigate dense settings. Our participants described care as a negotiated, family-mediated practice and highlighted the need for quick-use templates and audit logs that facilitate predictable and reversible sharing of information. Implementing per-metric toggles, role/situation templates (e.g., monitor/nudge/emergency (Binda et al., 2018)), and an exportable consent log addresses practical needs. Such controls also lower the social cost of sharing (users can hand off only what is needed to a spouse or clinician) and simplify onboarding for less technical family members. |
| R2 | Multi-mode prompting and stakeholder-specific escalation: Prompting must be layered depending on the escalation of the situation. While gentle, scheduled reminders, context-aware nudges, and configurable escalation paths are already available for the individual, it is essential to modify these features to extend them to chosen family members and, if necessary, to clinicians. Participants in our workshops described social prompts and clinician requests as the most effective triggers for action. A system that supports quiet hours for not just the user but all stakeholders involved in their care, snooze/acknowledge responses, and an explicit escalation chain both reduces nuisance alerts and creates clear pathways to care when thresholds persist. Empirical work (Orzikulova et al., 2024) on adaptive JITAI-style prompting and explainable timing supports the development of escalation policies that are both behaviorally effective and intelligible to users, and it is important to extend them to other stakeholders involved. |
| R3 | Privacy-preserving defaults and ethical safeguards: Systems must default to minimal sharing, offer clear and plain-language privacy summaries, provide straightforward retention controls (such as auto-delete after a specified number of days), and include revoke capabilities. Where PHI use carries social risk, governance by local ethics or community advisory boards should guide deployment. Our data show that although privacy concerns were limited and stakeholder-specific, trust concerns were a significant concern. Privacy-preserving defaults and visible, example-based consent dialogues reduce surprise and help users make informed choices (as observed through (Mackert et al., 2016; Meyers et al., 2020)). Recent research (Lee et al., 2024a) on privacy awareness tools and consent interfaces offers practical patterns for surfacing downstream uses and defaults. |
| R4 | Trusted-circle workflows to overcome adoption barriers: Because family networks were a primary adoption channel in our sample, systems should provide low-friction “care circles” with role-based permissions (ensuring sensitive information is shared with appropriate stakeholders while respecting the boundaries), simple invitation and verification flows, and time- or event-limited sharing. Participants valued the ability to set role templates that remove the need to micromanage every share. Embedding onboarding that explains each role’s responsibilities further reduces confusion and social friction. Prior work on family fitness (Li et al., 2021), intergenerational sharing (Binda et al., 2017; Cagiltay and Mutlu, 2024; Lee et al., 2024b), and women’s data ecosystems (Grimme et al., 2024; Karusala et al., 2023a) corroborates that role-sensitive sharing affordances increase both engagement and perceived safety. |
| R5 | Culturally-tailored motivation and contextual goals: Engagement mechanisms and goal framings should connect to locally meaningful activities, such as household tasks, festival periods, or family milestones, rather than focusing solely on gym-centric targets. Participants told us that culturally relevant framing and local language content make routine monitoring feel relevant and social. Embedding goal templates and coaching snippets that reference everyday practices (for example, festival-timed challenges or choreography-linked step goals) aligns PHI with lived routines and increases the likelihood that the system will be used as intended. Prior HCI work (Cha et al., 2025) on family and community-centered tracking supports localized, socially grounded motivators. |
| R6 | Clinician-ready summaries & one-click exports: Existing sharing mechanisms from management-type applications already provide a blueprint for promoting better cohesion among stakeholders. To translate PHI into clinical action, provide concise, one-page clinician summaries that highlight time-windowed trends, the most recent abnormal events, and provenance notes, along with export templates (PDF/SMART-on-FHIR/HL7 snippets) that cleanly map into clinical workflows. Participants explicitly framed such exports as enablers of earlier care. Making a single “clinician view” and export button available at the point of sharing reduces friction during visits and increases the likelihood clinicians will consider PHI. Existing implementations (Mandl et al., 2020) and SMART/HL7 guidance show feasible export patterns for embedding wearable summaries into care. |
@@ 5.2. Recommendations for Designers and Developers
| R7 | Supporting digital health literacy: Rather than exposing raw values and remaining black boxes (as observed in prior work (Shin et al., 2019; Zhang and Shahriar, 2020)), interfaces should translate detected patterns into concise, prescriptive “what to do” guidance, accompanied by optional clinical details, behind a progressive-disclosure affordance. Our respondents requested micro-guidance (e.g., “If X Y for Z days re-measure/share with clinician”) and examples that map metrics to everyday, embodied cues. Presenting single-sentence action steps upfront, along with a link to clinician-level notes, reduces the interpretation burden while preserving rigor for expert users. Prior literature on health literacy (Bagalkot et al., 2020; Su, 2023; Teixeira Lopes and Ramos, 2020; Egede et al., 2021; Teixeira Lopes and Henriques, 2024; Noblin et al., 2012; Kreps, 2017; Lee et al., 2022; Kemp et al., 2021) and rampant medical misinformation (Pasi, 2024) are critical indicators for the poor adoption of health technologies. HCI work (Wolf, 2019) on explainability and patient-facing explanations indicates that scenario-based, plain-language explanations materially improve comprehension and trust. Moreover, authoritative Indian health association articles rank poorly in search results from Indian IPs, often supplanted by vague content, whereas Western IP searches surface reliable sources prominently. This disparity underscores the need for authentic, reliable, and vetted user-friendly health information platforms, enabling Indians to access trustworthy guidance without hesitation. And since the pandemic has driven more health-conscious behaviors, implementing the long-overdue 2005 school-health curriculum recommendation (MoHFW, 2005) could foster health-conscious behavior from a young age (Kann et al., 2007; Pulimeno et al., 2020; Auld et al., 2020; Benes and Alperin, 2022; Sulz et al., 2024). |
| R8 | Tracking Provenance for reliability: Each derived metric should surface a compact provenance and confidence statement (for example, “Estimated from wrist PPG; confidence: medium; last sync 2h ago”), so users know how much to rely on a reading and whether to corroborate it with another device or a clinician. Study participants described cross-checking devices before acting. Lightweight badges and an optional provenance modal (utilizing sampling rate, sensor type, and last calibration, among other factors) reduce the cognitive friction of validation, helping users decide when to act versus when to re-measure. Technical and HCI literature (Cho et al., 2021) on wearable uncertainty and person-generated data quality and literature on user perception and trust (Huhn et al., 2022; Volpato et al., 2021; Kang and Exworthy, 2022; Singh et al., 2022a; Samhale, 2022; Zhang et al., 2022; Karnatak et al., 2023; Thapa et al., 2023) supports surfacing these dimensions to users, critical for holistic support. |
| R9 | Interoperability and data portability: Design the data layer around standard import/export formats (FHIR/CSV, OAuth connectors), source tagging, and deduplication/time-alignment so users and clinicians can assemble coherent longitudinal views from heterogeneous devices. Our participants framed ecosystem fragmentation as a barrier to meaningful histories. Supporting connector flows and clearly labelling provenance reduces the manual effort of assembling past data and enables portability across apps and EHRs. While prior work (Jones et al., 2012) indicates the IT infrastructure is still broken and fragmented, recent HCI and health-informatics work (Bathelt et al., 2022) emphasizes SMART-on-FHIR patterns for wearable integration as practical and interoperable. |
| R10 | Measure proactivity—evaluation & feedback loops: Finally, teams should instrument and report operational metrics that capture proactivity, such as the Proactivity Rate (the percentage of flagged events that result in contact within 7 days), Alert-to-Action latency, clinician-confirmed early detections, and other relevant metrics. Our participants linked specific features (family nudges, clinician summaries, confidence badges) to earlier action. Tracking these metrics in pilots and using them to pilot-test design choices produces the evidence needed to iterate responsibly and demonstrate clinical value. Work (Cho et al., 2021) on person-generated data quality and wearable evaluation offers validated measurement constructs that can seed these dashboards. |
Each recommendation is grounded in participant evidence and validated by the co-design prototype as a proof of concept (where relevant). This work contributes to PHI theory by translating ecological PHI (Murnane/Bronfenbrenner) into an operational PHI-for-proactivity map. Rather than a descriptive nesting of contexts alone, PHI-Proact prescribes concrete UI/UX and system levers (agency templates, confidence indicators, clinician export formats) that operationalize ecological constructs. In non-WEIRD settings, where family caregiving and systemic fragmentation are salient, this mapping demonstrates how ecological theory can directly inform design and evaluation to support proactive collective care, concretely answering our RQ2.
5.3. Limitations and Future Work
@@
While there are novel recommendations on personal, public, and cultural aspects of PHI adoption, the authors acknowledge several limitations in this study. First, our analysis draws on surveys, interviews, and co-design workshops; these methods elicit rich qualitative insights but are not a substitute for longitudinal field data, thereby limiting the ability to observe how these factors may evolve over time. Self-reported behaviour and intentions may not directly translate to sustained real-world behaviour change. As generalizability was not an objective of this work, the focus on urban, middle- and upper-middle-class participants restricts the application of the findings to other populations, where different adoption barriers may exist. Finally, the reliance on self-reported data also introduces potential bias, as participants may inaccurately report their use of PHI, utility, and adoption for proactive care. The prototype testing was qualitative and short-term, serving as a proof-of-concept for design directions rather than evidence of real-world impact. The prototype requires field deployment and longitudinal measurement to validate the effect of the recommended features on proactivity rate and clinical outcomes.
@@
Building upon our observations and the strategic direction outlined, we propose the design and development of a proof-of-concept application aimed at addressing general health, wellbeing, and medical inquiries. This system will also assist users in interpreting and contextualizing their personal health data aggregated from diverse sources, presenting the information in a format that is both comprehensible and actionable. Leveraging advancements in artificial intelligence, particularly finetuned open-source large language models (LLMs), the application will integrate data from clinical documents (e.g., physician reports, laboratory results, and medication records) alongside information from wearable health devices. This multimodal data synthesis will enable the generation of personalized, holistic insights (similar to the objective of Moore et al.’s work (Moore et al., 2022)), thereby enhancing user engagement and health awareness (achieving observations akin to Lin et al. (Lin et al., 2024)). The interface will employ a conversational, chatbot-style design to facilitate intuitive interaction, mediate family sharing norms, and surface literacy-appropriate explanations. Beyond helping users understand their PHI, the system will also provide access to evidence-based health, medical, and wellness content. By sourcing validated information from reputable sources, the application seeks to address the pervasive issue of health misinformation and contribute to the improvement of public health literacy.